CONFORMATION AND AGGREGATION OF FLUORINATED ALCOHOLS

\underline{TINA SCHARGE} and MARTIN A. SUHM, Institut für Physikalische Chemie, Universität Göttingen, Tammannstr. 6, 37077 Göttingen, Germany.

Replacement of hydrogen by fluorine in aliphatic alcohols leads to rich isomerism in the associated clusters, involving conformation, hydrogen bond topology and chiral recognition issues.^{a} Besides the classical OH⋯O hydrogen bond the aggregates are stabilized by two weaker interactions: OH⋯F and CH⋯F. The aggregates are formed in a jet expansion and characterized by FTIR spectroscopy, based on their stagnation pressure dependence and on argon coating effects. The experimental data are supported by quantum chemical calculations.

aT. Scharge, C. Emmeluth, T. Häber, M. A. Suhm, Competing hydrogen bond topologies in 2-fluoroethanol dimers, J. Mol.

Struct., (2006), in press.